OpenStudy (anonymous):

x-y=2 graphing using 2,3 for x

jimthompson5910 (jim_thompson5910):

To graph any line, all you need is two points. To find a point that lies on the line, plug in any x value to find its corresponding y value. So plug in x=2 to get 2-y = 2 and solve for y to get y = 0 So one point is (2,0) Now plug in x=3 to get 3-y = 2 and solve for y to get y = 1 So another point is (3,1) Now that you have these two points, plot them on a coordinate grid and draw a straight line through them to graph x - y = 2